I need to solve the3c+2d=2
d=4

Answer by mananth(16949)

3c+2d=2
d=4
..
plug value of d
3c+2*4=2
3c+8=2
add -8
3c+8-8=-8+2
3c=-6
divide by 3
3c/3=-6/3
c=-2
...